F1 = 40N, 30º SE; F2 = 25N, 40º NW; and F3 = 30N, 50º NE, what is the resultant vector R?


1
Expert's answer
2021-09-29T00:27:39-0400
"F_1=(40\sin(180\degree-30\degree), 40\cos(180\degree-30\degree))"

"F_2=(25\sin(-40\degree), 25\cos(-40\degree))"

"F_3=(30\sin(50\degree), 30\cos(50\degree))"

"R=F_1+F_2+F_3"

"=(40\sin(150\degree)-25\sin(40\degree)+30\sin(50\degree),"

"40\cos(150\degree)-\cos(40\degree)+30\cos(50\degree))"

"\approx(26.911643, -34.508499)"

"|R|=\sqrt{(26.911643)^2+(-34.508499)^2}"

"=43.76\ N"

"\tan \theta=\dfrac{-34.508499}{26.911643}"

"\theta=-52.05\degree"

"R = 43.76\ N, 37.95º SE"